A plate in the shape of a triangle is submerged vertically and the base which is 3 ft long is in the surface of the water. Find the altitude of the triangle if the force due to the water pressure against the face of the plate is 50w where w is the weight per cu. ft. of water.


1
Expert's answer
2021-05-03T10:32:50-0400

"S=\frac 12 lh,"

"p=50,"

"p=\frac FS,"

"\frac{W}{\frac 12 lh}=50,"

"h=\frac{W}{25l}=\frac{W}{75}~ft."